# DoraemonGPT : Toward Understanding Dynamic Scenes with Large Language Models (Exemplified as A Video Agent)

### v3

@@ Abstract
Recent LLM-driven visual agents mainly focus on solving image-based tasks, which limits their ability to understand dynamic scenes, making it far from real-life applications like guiding students in laboratory experiments and identifying their mistakes. Hence, this paper explores DoraemonGPT, a comprehensive and conceptually elegant system driven by LLMs to understand dynamic scenes. Considering the video modality better reflects the ever-changing nature of real-world scenarios, we exemplify DoraemonGPT as a video agent. Given a video with a question/task, DoraemonGPT begins by converting the input video into a symbolic memory that stores task-related attributes. This structured representation allows for spatial-temporal querying and reasoning by well-designed sub-task tools, resulting in concise intermediate results. Recognizing that LLMs have limited internal knowledge when it comes to specialized domains (e.g., analyzing the scientific principles underlying experiments), we incorporate plug-and-play tools to assess external knowledge and address tasks across different domains. Moreover, a novel LLM-driven planner based on Monte Carlo Tree Search is introduced to explore the large planning space for scheduling various tools. The planner iteratively finds feasible solutions by backpropagating the result’s reward, and multiple solutions can be summarized into an improved final answer. We extensively evaluate DoraemonGPT’s effectiveness on three benchmarks and several in-the-wild scenarios.
@@ 1 Introduction
Toward understanding dynamic scenes, developing LLM-driven agents to handle videos is of great significance yet involves grand challenges: i) Spatial-temporal Reasoning. Reasoning about the relationships of instances is crucial for task decomposing and decision-making. Such relationships may be relevant to space (Santoro et al., 2017), time (Zhou et al., 2018), or their spatial-temporal combination. i) Larger Planning Space. Compared to the image modality, high-level semantics about actions and their intentions can typically only be inferred from temporal visual observations (Tapaswi et al., 2016). In other words, the inference of temporal semantics is necessary and will enlarge the search space of decomposing dynamic video tasks. iii) Limited Internal Knowledge. LLMs can not encode all the knowledge required for understanding every video due to the ever-changing nature of the real world and/or the lack of learning on proprietary datasets (Peng et al., 2023a).
@@ 2 DoraemonGPT
| Category | YOLOv8 (Jocher et al., 2023)/Grounding DINO (Liu et al., 2023b) | The category of an instance, e.g., person | ||
@@
| Timestamp | - | The timestap of a frame/clip | ||
@@ 2.1 Task-related Symbolic Memory (TSM)
Time-dominant memory focuses on constructing temporal-related information of the video regarding the question. It requires comprehending the content throughout the video. The attributes stored in this memory include timestamp, audio content by ASR (Radford et al., 2023), optical content by OCR (PaddlePaddle, 2023), frame-level captioning by (Li et al., 2022, 2023a; Dai et al., 2023), clip-level captioning by dedupicating similar and continuous frame-level results, etc.
@@ 2.3 Monte Carlo Tree Search (MCTS) Planner
Previous LLM-driven planners (Shen et al., 2023; Surís et al., 2023; Gupta & Kembhavi, 2023) decompose the given into an action/sub-task sequence and solve it step by step. Such a strategy can be seen as a greedy search method which generates a chain of action nodes until the final answer. However, we consider the large planning space of solving dynamic video tasks as a tree, and a single attempt may not yield the correct result, or better solutions may exist. To efficiently explore the planning space, we propose a novel tree-search-like planner equipped with MCTS (Coulom, 2006; Kocsis & Szepesvári, 2006; Browne et al., 2012), which has shown its practicality in searching large trees.
@@
Reward Back-propagation. After obtaining a leaf/outcome node , we will gradually propagate its reward to its ancestor nodes until . Here we consider two kinds of reward:
@@
For simplicity, let be a positive base reward, we set for Failure and Non-failure, respectively. According to (Liu et al., 2023a), the outcome produced by LLMs is more related to the context at the beginning (the initial prompts) and the end (the final nodes). We consider that more reward should be applied to nodes close to . Thus, the backpropagation function is formulated as , where denotes the node distance between and , and is a hyper-parameter for controlling the decay rate of the reward. The further the node distance, the greater the reward decay ratio, .
@@ 3.1 Experimental Setup
Implementation Details. We use GPT-3.5-turbo API provided by OpenAI as our LLM. As summarized in Table 1, we use BLIP series (Dai et al., 2023) for captioning, YOLOv8 (Jocher et al., 2023) and Deep OC-Sort (Maggiolino et al., 2023) for object tracking, PaddleOCR (PaddlePaddle, 2023) for OCR, InternVideo (Wang et al., 2022) for action recognition, and Whisper (Radford et al., 2023) for ASR. Our experiments are conducted under the in-context learning (ICL) setting.
@@
Reproducibility. Our algorithm is implemented in PyTorch and LangChain and tested on an NVIDIA Tesla A100 GPU. Code will be released to guarantee reproducibility.
@@ 3.2 Zero-shot Video Question Answering
3.3 Zero-shot Referring Object Segmentation
Ref-YouTube-VOS. We further evaluate DoraemonGPT against state-of-the-art supervised referring video object segmentation models and LLM-driven agents, as summarized in Table 2(b). Benefiting from our task-related symbolic memory, the DoraemonGPT (without learning on Ref-YouTube-VOS) can effectively ground video instances with textual description and remarkably surpasses recent supervised models (65.9% v.s. 64.8%), e.g., OnlineRefer (Wu et al., 2023b). By contrast, the agent competitor, ViperGPT, only achieves 26.6% due to a lack of well-designed video information memory, failing to ground the referred object or track it accurately in the video. Our superiority in referring video object segmentation further demonstrates the necessity of building a symbolic video memory. Please refer to A.4 for more qualitative results.
3.4 In-the-wild Example
@@ 3.5 Diagnostic Experiment
Multiple Solutions by MCTS Planner. We next study the influence of the number of answer candidates during the exploring process of our MCTS planner. When 1, the planner is equivalent to a greedy search, explores only a chain of nodes, and returns a single answer – the first node in LLM’s thinking that can be terminated without further exploration. As shown in Table 3(b), gradually increasing from 1 to 4 leads to better performance (i.e., 43.3 65.7). This supports our hypothesis that one single answer is far from enough to handle the larger planning space for dynamic modalities and proves the efficacy of our MCTS planner. Since the questions in NExT-QA (Xiao et al., 2021) are single-choice, exploring more answers does not always result in positive returns. We stop using 5 as the required number of API calls exceeds our budget.
Back-propogation in MCTS Planner. Then, we ablate the effect of the base reward and decay rate ( §2.3) that controls the exploring procedure of our MCTS planner. The results in Table 3(c) are stable regardless of the combination of and used. Hence, we set the slightly better one, and , as our default setting. We leave some special combinations in the next part (e.g., our MCTS planner becomes the depth-first search (DFS) when setting and 1 for both Failure and Non-failure cases).
Exploring Strategies used by Planner. Last, to verify the advantage of our MCTS planner, we compare MCTS with several standard exploring strategies, i.e., depth-first search (DFS), Root, which always selects the root node, and Uniform, which samples a node with equal probability. As shown in Table 3(d), we observe that their performance is suboptimal due to the inability to leverage the value/reward of the outcome leaf nodes and accordingly adjust their searching strategy. Compared to these naïve strategies, our MCTS planner adaptively samples a node with the guidance of reward back-propagation, which is more effective in a large solution space. These results further validate the superiority of the proposed MCTS planner.
@@ 4 Related Work
Multimodal Understanding. Before the emergence of LLMs, various efforts were made to create multimodal systems tailored for specific tasks (Anderson et al., 2018; Lu et al., 2019; Marino et al., 2019, 2021; Bain et al., 2021; Lei et al., 2021; Grauman et al., 2022; Lin et al., 2022; Li et al., 2023c; Wong et al., 2022; Chen et al., 2023a; Hu et al., 2020; Yang et al., 2021b). While these systems showed impressive performance in their respective domains, their applicability to broader, real-world scenarios was limited due to the lack of generalizability. Recent years have witnessed remarkable progress in general multimodal systems, due to the fast evolution of data volumes and computational resources. Specifically, Frozen (Tsimpoukelli et al., 2021) is a typical example; it demenstrates a feasible way to empower LLMs the ability to handle visual inputs. In the past few years, numerous efforts have been devoted to build large multimodal models (OpenAI, 2023; Driess et al., 2023; Zhu et al., 2023a; Li et al., 2022). Considering the training cost, several attempts (Li et al., 2023a; Yu et al., 2023) try to build zero-shot systems for various tasks. An alternative strategy (Xi et al., 2023), which will be detailed latter, involves combining multiple models or APIs to tackle compositional multimodal reasoning tasks. Our DoraemonGPT shares a similar spirit of decomposing complex tasks into simpler ones, but it is designed to solve complicated tasks for dynamic modalities in the real-world senarios.
@@
LLM-driven Modular Systems. Deconstructing complex tasks and merging the results from multiple intermediate steps is an innate human ability that drives the scientific and industrial communities (Newell et al., 1972; Wang & Chiew, 2010). Benefiting from the impressive emergent capabilities of LLMs, VisProg (Gupta & Kembhavi, 2023) pioneers the idea of addressing complex vision tasks through the decomposition of questions into manageable subtasks. Along this line, tremendous progress has been achieved, which can be divided into two categories according to reasoning styles: i) Reasoning with fixed paths (Gupta & Kembhavi, 2023; Wu et al., 2023a; Surís et al., 2023; Lu et al., 2023; Shen et al., 2023; Liang et al., 2023). They transform the given task into a ordered sequence of subtasks, each addressed by a specific module. For example, ViperGPT (Surís et al., 2023) treats the solving process as a Python program with manually designed APIs. Similarly, HuggingGPT (Shen et al., 2023) models task dependencies between numerous foundation models. ii) Reasoning with dynamic paths (Nakano et al., 2021; Schick et al., 2023; Yao et al., 2022; Yang et al., 2023; Gao et al., 2023a). Considering the intermediate results may not meet expectations, a promising avenue is to perform planning and executing concurrently. Such an interactive paradigm (Yao et al., 2022) provides a flexible and error-tolerant manner compared to those with fixed paths. Additionally, there are many agents focussing on other domains, e.g., planning in the open-world environments (Wang et al., 2023c; Yuan et al., 2023; Park et al., 2023), tool usage (Ruan et al., 2023; Qin et al., 2023), reinforcement learning (Shinn et al., 2023; Xu et al., 2023). This work focuses on computer vision only.
@@
Despite impressive, existing LLM-driven modular systems mainly focus on developing specific strategies to solve composition tasks for static modalities, ignoring the fundamental gaps between static and dynamic modalities, which is a pivotal aspect towards achieving artificial general intelligence (AGI) (Goertzel, 2014). These works, to some extent, can be seen as a subset of our system. Though exceptions exist (Surís et al., 2023; Li et al., 2023b; Wang et al., 2023a; Gao et al., 2023a), in general they are scattered, lacking systematic study, e.g., simply treating the video as a sequence of images (Surís et al., 2023) or building a chatbot based on pre-extracted information of the given video (Li et al., 2023b; Wang et al., 2023a). In sharp contrast, we take the video and task as a whole, resulting in a compact, task-related memory. The reasoning paths of our system are dynamic powered by the MCTS planner. In addition to facilitate the answer searching, MCTS planner has the potential to find multiple possible candidates. This is crucial for questions with open-ended answer.
@@
LLMs with External Memory. How to effectively design prompt templates, known as prompt engnieering, is of importance for accurate LLM responses (Zhou et al., 2022; Wang et al., 2023b). One of the areas in the spotlight is memory-augmented LLMs (Wu et al., 2022c, b; Zhong et al., 2022; Lewis et al., 2020; Guu et al., 2020; Izacard et al., 2022; Khattab et al., 2022; Park et al., 2023; Cheng et al., 2022; Sun et al., 2023; Hu et al., 2023). In general, training-free memories can be divided into: i) Textual memory (Zhu et al., 2023b; Park et al., 2023). In this kind of memory, the long contexts LLMs cannot handle (e.g., books) are stored as embeddings, which can be further retrieved by computing similarity. A typical example is the document question answering shown in LangChain111https://docs.langchain.com/docs/use-cases/qa-docs. ii) Symbolic memory. It models memory as structured represtations with corresponding symbolic languages, e.g., codes for programing language (Cheng et al., 2022), execution commands for Excel222https://chatexcel.com/, and structured query language (SQL) for databases (Sun et al., 2023; Hu et al., 2023). Different from techniques (Dao et al., 2022; Dao, 2023; Ratner et al., 2023; Hao et al., 2022; Chen et al., 2023b; Mu et al., 2023; Mohtashami & Jaggi, 2023; Peng et al., 2023b) that directly extend the context window of LLMs, memory-augmented LLMs use retrieval-based approaches to bypass the limitation of context length. This is more favored because: (i) it is a plug-and-play module without any fine-tuning or architectural modifications; and (ii) concurrent works (Shi et al., 2023; Liu et al., 2023a) suggest that LLMs may be distracted or lost while encounting irrelevant or long contexts. By absorbing their ideas of memory organization, we construct a request-related database, which stores intance-aware and instance-agnostic information into indivisual tables. To retrieve the relevant information, we explicitly define several sub-task tools based on prompt templates and SQL, with respect to different purposes. With a broader view, our multi-source knowledge, which is a complementary module to provide reliable guidances in specific domains, can also be considered as a hybrid of external memories.

### v4

@@ Abstract
Recent LLM-driven visual agents mainly focus on solving image-based tasks, which limits their ability to understand dynamic scenes, making it far from real-life applications like guiding students in laboratory experiments and identifying their mistakes. Hence, this paper explores DoraemonGPT, a comprehensive and conceptually elegant system driven by LLMs to understand dynamic scenes. Considering the video modality better reflects the ever-changing nature of real-world scenarios, we exemplify DoraemonGPT as a video agent. Given a video with a question/task, DoraemonGPT begins by converting the input video into a symbolic memory that stores task-related attributes. This structured representation allows for spatial-temporal querying and reasoning by well-designed sub-task tools, resulting in concise intermediate results. Recognizing that LLMs have limited internal knowledge when it comes to specialized domains (e.g., analyzing the scientific principles underlying experiments), we incorporate plug-and-play tools to assess external knowledge and address tasks across different domains. Moreover, a novel LLM-driven planner based on Monte Carlo Tree Search is introduced to explore the large planning space for scheduling various tools. The planner iteratively finds feasible solutions by backpropagating the result’s reward, and multiple solutions can be summarized into an improved final answer. We extensively evaluate DoraemonGPT’s effectiveness on three benchmarks and several in-the-wild scenarios. Project page: https://z-x-yang.github.io/doraemon-gpt.
@@ 1 Introduction
Toward understanding dynamic scenes, developing LLM-driven agents to handle videos is of great significance yet involves grand challenges: i) Spatial-temporal Reasoning. Reasoning about the relationships of instances is crucial for task decomposing and decision-making. Such relationships may be relevant to space (Santoro et al., 2017), time (Zhou et al., 2018), or their spatial-temporal combination. ii) Larger Planning Space. Compared to the image modality, high-level semantics about actions and their intentions can typically only be inferred from temporal visual observations (Tapaswi et al., 2016). In other words, the inference of temporal semantics is necessary and will enlarge the search space of decomposing dynamic video tasks. iii) Limited Internal Knowledge. LLMs can not encode all the knowledge required for understanding every video due to the ever-changing nature of the real world and/or the lack of learning on proprietary datasets (Peng et al., 2023a).
@@ 2 DoraemonGPT
| Category | YOLOv8 (Jocher et al., 2023)/Grounding DINO (Liu et al., 2023c) | The category of an instance, e.g., person | ||
@@
| Timestamp | - | The timestamp of a frame/clip | ||
@@ 2.1 Task-related Symbolic Memory (TSM)
Time-dominant memory focuses on constructing temporal-related information of the video. It requires comprehending the content throughout the video. The attributes stored in this memory include timestamp, audio content by ASR (Radford et al., 2023), optical content by OCR (PaddlePaddle, 2023), frame-level captioning by BLIPs (Li et al., 2022, 2023a; Dai et al., 2023), clip-level captioning by dedupicating similar and continuous frame-level results, etc.
@@ 2.3 Monte Carlo Tree Search (MCTS) Planner
Previous LLM-driven planners (Shen et al., 2023; Surís et al., 2023; Gupta & Kembhavi, 2023) decompose the given into an action/sub-task sequence and solve it step by step. Such a strategy can be seen as a greedy search method that generates a chain of action nodes until the final answer. Similar to some works (Yao et al., 2024; Zhuang et al., 2024), we consider the large planning space of LLM-driven planning as a tree. Moreover, we consider that a single attempt may not yield the correct result or that better solutions may exist. To efficiently explore the planning space, we propose a novel tree-search-like planner equipped with MCTS (Coulom, 2006; Kocsis & Szepesvári, 2006; Browne et al., 2012), which is practical in searching large trees.
@@
Reward Back-propagation. After obtaining a leaf/outcome node , we will gradually propagate its reward to its ancestor nodes until . Here, we consider two kinds of reward:
@@
For simplicity, let be a positive base reward, we set for Failure and Non-failure, respectively. According to (Liu et al., 2023a), the outcome produced by LLMs is more related to the context at the beginning (the initial prompts) and the end (the final nodes). We believe more rewards should be applied to nodes close to . Thus, the backpropagation function is formulated as , where denotes the node distance between and , and is a hyper-parameter for controlling the decay rate of the reward. The further the node distance, the greater the reward decay ratio, . Commonly, setting a higher / will increase the probability of the expanded node approaching the leaf node (of a non-failure answer).
@@ 3.1 Experimental Setup
Implementation Details. We use GPT-3.5-turbo API provided by OpenAI as our LLM. As summarized in Table 1, we use BLIP series (Dai et al., 2023) for captioning, YOLOv8 (Jocher et al., 2023) and Deep OC-Sort (Maggiolino et al., 2023) for object tracking, PaddleOCR (PaddlePaddle, 2023) for OCR, InternVideo (Wang et al., 2022) for action recognition, and Whisper (Radford et al., 2023) for ASR. Our experiments are conducted under the in-context learning (ICL) setting. The external knowledge is not used in quantitative or qualitative comparisons to ensure fairness.
@@ 3.2 Zero-shot Video Question Answering
3.3 Zero-shot Referring Object Segmentation
Ref-YouTube-VOS. We further evaluate DoraemonGPT against state-of-the-art supervised referring video object segmentation models and LLM-driven agents, as summarized in Table 2(b). Benefiting from our task-related symbolic memory, the DoraemonGPT (without learning on Ref-YouTube-VOS) can effectively ground video instances with textual description and remarkably surpasses recent supervised models (65.9% v.s. 64.8%), e.g., OnlineRefer (Wu et al., 2023b). By contrast, the agent competitor, ViperGPT, only achieves 26.6% due to a lack of well-designed video information memory, failing to ground the referred object or track it accurately in the video.
As depicted in Fig. 4, our DoraemonGPT demonstrates a higher level of accuracy in identifying, tracking, and segmenting the referred objects mentioned by the users. In contrast, ViperGPT encounters failure cases where the recognized objects do not completely match the semantic and descriptive aspects. Our superiority in referring video object segmentation further demonstrates the necessity of building a symbolic video memory.
3.4 In-the-wild Example
@@ 3.5 Diagnostic Experiment
Multiple Solutions by MCTS Planner. We will next study the influence of the number of answer candidates during the exploration process of our MCTS planner. When 1, the planner is equivalent to a greedy search, explores only a chain of nodes, and returns a single answer – the first node in LLM’s thinking that can be terminated without further exploration. As shown in Table 3(c), gradually increasing from 1 to 4 leads to better performance (i.e., 43.3 65.7). This supports our hypothesis that one single answer is far from enough to handle the larger planning space for dynamic modalities and proves the efficacy of our MCTS planner. Since the questions in NExT-QA (Xiao et al., 2021) are single-choice, exploring more answers does not always result in positive returns. We stop using 5 as the required number of API calls exceeds our budget.
Back-propogation in MCTS Planner. Then, we ablate the effect of the base reward and decay rate ( §2.3) that controls the exploring procedure of our MCTS planner. The results in Table 3(d) are stable regardless of the combination of and used. Hence, we set the slightly better one, and , as our default setting. We leave some special combinations in the next part (e.g., our MCTS planner becomes the depth-first search (DFS) when setting and 1 for both Failure and Non-failure cases).
Exploring Strategies used by Planner. Last, to verify the advantage of our MCTS planner, we compare MCTS with several standard exploring strategies, i.e., depth-first search (DFS), Root, which always selects the root node, and Uniform, which samples a node with equal probability. As shown in Table 3(e), we observe that their performance is suboptimal due to the inability to leverage the value/reward of the outcome leaf nodes and accordingly adjust their searching strategy. Compared to these naïve strategies, our MCTS planner adaptively samples a node with the guidance of reward back-propagation, which is more effective in a large solution space. These results further validate the superiority of the proposed MCTS planner.
Additionally, compared to Uniform, DFS notably exhibits superior performance on temporal questions yet demonstrates comparable or even inferior performance on descriptive and causal questions. We hypothesize that this discrepancy arises because temporal questions often include cues (e.g., at the beginning) indicating the specific period within the video where the correct answer can be found. DFS can exploit these cues, whereas Uniform cannot.
Impact of Captioning Models. We further conduct experiments regarding one of the most important models, i.e., BLIP series, which is the essential tool for DoraemonGPT to perceive and recognize visual input. As seen in Table 3(b), the model with instruction tuning, i.e., InstructBLIP, demonstrates better performance. This suggests that DoraemonGPT can benefit from the development of stronger foundation models. More discussion of foundation models’ impact on DoraemonGPT is supplied in A.4.
@@ 4 Related Work
Multimodal Understanding. Various efforts were made to create multimodal systems tailored for specific tasks (Lu et al., 2019; Marino et al., 2019, 2021; Bain et al., 2021; Lei et al., 2021; Grauman et al., 2022; Lin et al., 2022; Li et al., 2023d, c; Lei et al., 2024; Wong et al., 2022; Chen et al., 2023a; Hu et al., 2020; Yang et al., 2021b). While these systems showed impressive performance in their respective domains, their applicability to broader, real-world scenarios was limited due to the lack of generalizability. Recent years have witnessed remarkable progress in general multimodal systems, due to the fast evolution of data volumes and computational resources. Specifically, Frozen (Tsimpoukelli et al., 2021) is a typical example; it demonstrates a feasible way to empower LLMs with the ability to handle visual inputs. In the past few years, numerous efforts have been devoted to building large multimodal models (OpenAI, 2023; Driess et al., 2023; Zhu et al., 2023a; Li et al., 2022). Considering the training cost, several attempts (Li et al., 2023a; Yu et al., 2023) try to build zero-shot systems for various tasks. An alternative strategy (Xi et al., 2023), which will be detailed later, involves combining multiple models or APIs to tackle compositional multimodal reasoning tasks. Our DoraemonGPT shares a similar spirit of decomposing complex tasks into simpler ones, but it is designed to solve complicated tasks for dynamic modalities in real-world scenarios.
@@
LLM-driven Modular Systems. Deconstructing complex tasks and merging the results from multiple intermediate steps is an innate human ability that drives the scientific and industrial communities (Newell et al., 1972; Wang & Chiew, 2010). Benefiting from the impressive emergent capabilities of LLMs, VisProg (Gupta & Kembhavi, 2023) pioneers the idea of addressing complex vision tasks through the decomposition of questions into manageable subtasks. Along this line, tremendous progress has been achieved, which can be divided into two categories according to reasoning styles: i) Reasoning with fixed paths (Gupta & Kembhavi, 2023; Wu et al., 2023a; Surís et al., 2023; Lu et al., 2023; Shen et al., 2023; Liang et al., 2023). They transform the given task into an ordered sequence of subtasks, each addressed by a specific module. For example, ViperGPT (Surís et al., 2023) treats the solving process as a Python program with manually designed APIs. Similarly, HuggingGPT (Shen et al., 2023) models task dependencies between numerous foundation models. ii) Reasoning with dynamic paths (Nakano et al., 2021; Schick et al., 2023; Yao et al., 2022; Yang et al., 2023; Gao et al., 2023a). Considering the intermediate results may not meet expectations, a promising avenue is to perform planning and executing concurrently. Such an interactive paradigm (Yao et al., 2022) provides a flexible and error-tolerant manner compared to those with fixed paths. Additionally, there are many agents focussing on other domains, e.g., planning in the open-world environments (Wang et al., 2023c; Yuan et al., 2023; Park et al., 2023), tool usage (Ruan et al., 2023; Qin et al., 2023), reinforcement learning (Shinn et al., 2023; Xu et al., 2023). This work focuses on computer vision only.
@@
Despite impressive, existing LLM-driven modular systems mainly focus on developing specific strategies to solve composition tasks for static modalities, ignoring the fundamental gaps between static and dynamic modalities, which is a pivotal aspect towards achieving artificial general intelligence (AGI) (Goertzel, 2014). These works, to some extent, can be seen as a subset of our system. Though exceptions exist (Surís et al., 2023; Li et al., 2023b; Wang et al., 2023a; Gao et al., 2023a), in general, they are scattered, lacking systematic study, e.g., simply treating the video as a sequence of images (Surís et al., 2023) or building a chatbot based on pre-extracted information of the given video (Li et al., 2023b; Wang et al., 2023a). In sharp contrast, we take the video and task as a whole, resulting in a compact, task-related memory. The reasoning paths of our system are powered by the MCTS planner. In addition to facilitating the answer searching, the MCTS planner has the potential to find multiple possible candidates. This is crucial for questions with open-ended answers.
@@
LLMs with External Memory. How to effectively design prompt templates, known as prompt engineering, is of importance for accurate LLM responses (Zhou et al., 2022; Wang et al., 2023b). One of the areas in the spotlight is memory-augmented LLMs (Wu et al., 2022c, b; Zhong et al., 2022; Lewis et al., 2020; Guu et al., 2020; Izacard et al., 2022; Khattab et al., 2022; Park et al., 2023; Cheng et al., 2022; Sun et al., 2023; Hu et al., 2023). In general, training-free memories can be divided into: i) Textual memory (Zhu et al., 2023b; Park et al., 2023). In this kind of memory, the long contexts LLMs cannot handle (e.g., books) are stored as embeddings, which can be further retrieved by computing similarity. A typical example is the document question answering shown in LangChain111https://docs.langchain.com/docs/use-cases/qa-docs. ii) Symbolic memory. It models memory as structured representations with corresponding symbolic languages, e.g., codes for programming language (Cheng et al., 2022), execution commands for Excel222https://chatexcel.com/, and structured query language (SQL) for databases (Sun et al., 2023; Hu et al., 2023). Different from techniques (Dao et al., 2022; Dao, 2023; Ratner et al., 2023; Hao et al., 2022; Chen et al., 2023b; Mu et al., 2023; Mohtashami & Jaggi, 2023; Peng et al., 2023b) that directly extend the context window of LLMs, memory-augmented LLMs use retrieval-based approaches to bypass the limitation of context length. This is more favored because (i) it is a plug-and-play module without any fine-tuning or architectural modifications, and (ii) concurrent works (Shi et al., 2023; Liu et al., 2023a) suggest that LLMs may be distracted or lost while encountering irrelevant or long contexts. By absorbing their ideas of memory organization, we construct a request-related database, which stores instance-aware and instance-agnostic information in individual tables. To retrieve the relevant information, we explicitly define several sub-task tools based on prompt templates and SQL, with respect to different purposes. With a broader view, our multi-source knowledge, which is a complementary module to provide reliable guidance in specific domains, can also be considered as a hybrid of external memories.